OpenStudy (anonymous):

Help please. Cant figure it out. A plumber charges 52.50 per hour plus $75.00 service charge. Your fathers firm hires him to fix some leaky pipes. Find the total charges if it takes the plumber 4.5 hours to complete the task.

OpenStudy (anonymous):

well, the plumber charges 52.50 per hour plus a service charge of 75. so what you are going to do is multiply 52.50 times 4.5 to get the total cost of the hours worked, then just add the $75 service charge and you have the answer! :) 52.50(4.5) + 75= Total money owed :)

OpenStudy (anonymous):

311.25? Total money owned?

OpenStudy (anonymous):

Since, plumber charges 52.5/ hour therefore, charge for five hours = 5 * 52.50 =262.50 Dollars Total charge of the plumber = charge for five hours + service charge = 262.50 + 75 = $337.50 @Dreams100